# Gravitational Waves from Magnetars

### v1

@@ Abstract
We study the emission of gravitational waves produced by the magnetosphere of magnetars. We argue that several features in the spectrum could facilitate the identification of that source. In addition, in cases of extremely large magnetic fields we demonstrate that this emission can make the braking index of such stars to be well over 3, which is the standard prediction of the magnetic dipole radiation and aligned rotator mechanisms. A similar picture arises if one focuses on the second braking index. Moreover the braking index depends on both the rotational frequency and the strength of the magnetic field in striking difference from the other mechanisms. We also show that gravitational waves can be produced by polar gap regions due to their rapid charge-discharge process that takes place in timescales from nanosec to microsec. This can provide an alternative way to probe magnetars and test the polar gap model.
@@
Since the first detection of gravitational waves (GW) a few years ago LIGOScientific:2016aoc , a lot of effort has been concentrated on this front both theoretically and experimentally. With the advent of Pulsar Timing Arrays and a series of new interferometers that have been proposed and expected to enter operation in the near or far future, there has been a lot of interest for new astrophysical GW sources that can potentially reveal a lot of information regarding astrophysical and cosmological processes. One field that has been studied extensively for obvious reasons is the production of GW by neutron stars (NS). Due to their compactness, there is a lot of opportunity to emit GW under certain conditions. NS can get involved in binaries among themselves or with black holes. This is the strongest type of signal that can come from a NS and this has already been detected LIGOScientific:2017vwq . Other possible cases of GW emission by NS is via several non-radial oscillations modes that can be excited in principle either by internal processes or by encounters with bypassing objects. A rotating NS is more oblate in the equator. This by itself cannot produce any GW, since the latter requires a time varying quadrupole (or higher) moment. Therefore in case for example a bypassing object disturbs the rotation of the NS, a nonzero second derivative of the quadrupole moment can be created, leading to emission of GW. For gravitational waves produced by pulsating NS, see e.g. Andersson:2000mf .
@@
However, there is another subtle mechanism where GW can be produced by NS involving neither some internal process nor a bypassing object. As it is well known since the 60’s, a rotating NS with a magnetic field cannot be without the presence of a magnetosphere GJ . In case where the rotation and magnetic axes do not coincide, the co-rotating magnetosphere can create GW that in principle can be detected in future experiments. This scenario has been already studied in Nazari and Contopoulos where the GW produced by the magnetosphere were deduced. The focus of these two papers was on the detectability of these GW in upcoming interferometers. Continuous GW produced in the context of glitches has been considered in Yim:2023nda .
@@
In this paper we verify that indeed GW produced by the magnetosphere of NS with strong magnetic fields are within reach in upcoming experiments. In addition we show that this is a source of continuous GW with a spectrum consisting of two frequencies, i.e., the frequency of the NS rotation $f_{0}$ and $2f_{0}$ with a very specific relative amplitude between the two. Furthermore we focus on another aspect of this scenario. We study how the emission of such GW by the magnetosphere affects the first and second braking index of the NS. We shall argue that in case of magnetars with extremely high magnetic fields, the braking index can be substantially over 3 which is the number expected in standard scenarios where the NS deceleration takes place due to magnetic dipole radiation and/or via torques of escaping to infinity currents from the poles of the NS. We find that unlike the usual magnetic dipole and aligned rotator models where the braking index is independent of both the rotation frequency and the magnetic field, in high magnetic field magnetars, the braking index depends on both mentioned quantities. Similar conclusions can be drawn for the second braking index. We will show that small deviations from the standard values result from NS with fields of at least $10^{16}$ Gauss and significant deviations are expected for fields of $10^{18}$ to $10^{20}$ Gauss. Although such extremely high magnetic fields have not been observed in magnetars so far, this is not unexpected. Arguments based on the virial theorem suggest fields up to $10^{18}$ Gauss for NS with nuclear cores ShapiroLai , while for magnetars with a quark core this value can rise to $10^{20}$ Gauss Ferrer:2010wz . Hybrid stars with quark and nuclear matter and high magnetic fields have been studied in Sotani:2014rva . The effects of extremely high magnetic fields on the QCD equation of state (relevant also for NS) has been recently readdressed Fraga:2023cef ; Fraga:2023lzn .
@@ I Gravitational Waves produced by the Magnetosphere
$3\times 10^{-24}\left(\frac{B_{0}}{10^{15}\text{Gauss}}\right)^{2}\left(\frac{R}{12~{}\text{km}}\right)^{5}\left(\frac{\text{msec}}{P}\right)^{2}\left(\frac{\text{kpc}}{r}\right),$
@@
Figure 2: First (upper panels) and second (lower panels) braking indices for magnetars of different magnetic field as a function of Log P ($P$ being the period of the star in sec). Upper Left panel: From higher index value to lower, we have stars with $B=10^{20}$, $10^{19}$ and $10^{18.5}$ Gauss. Right panel: The same thing for $B=10^{17.5}$, $10^{17}$ and $10^{16.5}$ Gauss. Lower panels: Same as the upper panels for the second braking index. In all plots we use $R=20$ km and $\alpha=\pi/2$.
@@ II Braking Index
For any mechanism where $\dot{\omega}=-k\omega^{n}$ with $k$ independent of time, the braking index is n and the second braking index is $m=n(2n-1)$. By inspection of Eq. (13) both the aligned rotator model and the magnetic dipole model give $n=3$ and $m=15$. Note also that GW emission provides first and second braking indices $n=5$ and $m=45$ respectively. The co-existence of the electromagnetic emission mechanisms with the GW emission one will make both $n$ and $m$ depart from their corresponding values of 3 and 15. This is depicted in Fig. (2) where both the first and second braking indices are plotted as a function of the rotational period. Pulsars with magnetic fields as low as $\sim 10^{16}$ Gauss could cause deviations from $n=3$ starting from $10^{-4}$. Such accuracy can be anticipated for some pulsars in the future given the current one which can be as low as $10^{-3}$. However even larger magnetic fields of the order of $10^{18}$ Gauss can cause a significant change in both $n$ and $m$ as it can be seen. As expected, faster rotations cause in general larger deviations from the standard $n=3$ and $m=15$ values. Note however that this is not completely true when we look at potential pulsars with periods below 10 msec. As it can be seen in the figure, both $n$ and $m$ in this mechanism are not decreasing functions of the period within the whole range of parameters. Both indices increase in the msec range as the period increases reaching a maximum and after that indeed $n$ and $m$ become decreasing functions of the period. This is in distinct contrast to what happens to the braking index that is dominated by GW emission under the condition that the GW emission is due to misalignment between the principal axes and the rotation one, which is expected to have braking indices monotonically decreasing with the period. This feature is attributed to the nontrivial way that powers of $\omega R$ enter the expression of the quadrupole moment. It reflects the nontrivial way that the GW emission depends on the rotation and magnetic field. Ordinarily, higher frequency means larger GW amplitude. However in this case, higher frequency leads also to smaller light cylinder and thus smaller volume that contributes to the GW emission. Note here that for a large part of the studied parameter space, it appears that the deviation from the standard value of $m$ is roughly an order of magnitude larger than that of $n$. This is anticipated and certainly true when the deviations are small. Another characteristic feature which is distinctly different from the typical GW emission due to the misalignment of an oblate principal axis with respect to the rotation axis or due to the existence of bumps in the surface of the star is the fact that in our studied scenario, both indices depend strongly on the magnetic field unlike the former cases. The magnetic field of NS is typically estimated by the deceleration rate of the rotation. However in principle, a different independent estimate of the magnetic field could actually test our scenario. In any case, even if the magnetic field cannot be determined independently, knowledge of both $n$ and $m$ can potentially distinguish this deceleration mechanism from the standard one that is independent of the magnetic field since in our studied scenario both $n$ and $m$ depend on the magnetic field nontrivially. As a final comment it is worth looking at models of decaying magnetic field. In such a context the magnetic field can be correlated to the rotational frequency. This could provide another way to test this mechanism since it could make a concrete prediction on both indices once the correlation between the magnetic field and rotational frequency is established.
@@ III Gravitational Waves from the Polar Cap
where $z$ denotes the coordinate along the rotation axis. Note that during the charging of the gap $t=z-R-T/2$ for $-T/2\leq t\leq-T/2+h$. $T$ is the total period of the cycle charge-discharge that we will estimate later on. Within our approximation after the electric field reaches its maximum, it will stay there up to time $t=T/2-h$ where the spark ignites. After that moment the charge redistributes and reaches the Goldreich-Julian value that screens completely the electric field. Therefore after $t=T/2-h$ and until $T=T/2$, we have $t=T/2+R-z$. Note also that $z$ is measured from the center of mass which in good approximation is taken to be the center of the star. Fig. (3) illustrates the geometry of the polar cap region.
@@
By inspection of Eq. (20), for relatively slow rotations $\ddot{M}_{33}>>\ddot{M}_{11}$ because $R>>r_{p}\simeq R\sqrt{\omega R}$. To give an approximate estimate for the GW amplitude, we note that $g(t)$ is of order one for time between $-T/2\leq t\leq-T/2+h$ and $T/2-h\leq t\leq T/2$. As we mentioned earlier the Fourier analysis gives a peak at $f_{1}=h^{-1}$ and a smaller amplitude at $f_{0}=T^{-1}$, consistent with our argument here.
@@
[…]
@@
As we mentioned $\ddot{M}_{11}<<\ddot{M}_{33}$. Ignoring the factor $\sin^{2}\psi$ and setting the distance to the source to 1kpc, we can estimate the amplitude for different benchmark points presented in Table 1. $f_{0}$ and $f_{1}$ are the two frequencies mentioned earlier. The amplitude for $h_{+}$ that is presented refers to $f_{1}$. Depending on the exact form of the functions $f(t)$ or $g(t)$, the amplitude at $f_{0}$ will be somewhat suppressed compared to that at $f_{1}$ that we quote at the table. For all benchmark points we used a period for the magnetar of $10^{-3}$ sec. Although for the moment there are no current experiments probing the MHz-GHz range of frequencies, several proposals for future experiments exist, aiming exactly at detection of high frequency GW Aggarwal:2020olq . Table 1 of Aggarwal:2020olq gives a summary of all the high GW frequency proposals. Among the different proposals, detectors based on enhanced magnetic conversion could in principle operate at $\sim 10$ GHz with a sensitivity for the strain $10^{-30}$ to $10^{-26}$. Similarly for superconducting rings, the projected sensitivity at 10 GHz is $10^{-31}$. By inspection of Table 1 magnetic fields as low as $10^{14}$ Gauss are above the threshold in the GHz frequency range of the aforementioned proposals.

### v2

@@ Abstract
We study the emission of gravitational waves produced by the magnetosphere of magnetars. We argue that several features in the spectrum could facilitate the identification of that source. In addition, in cases of extremely large magnetic fields we demonstrate that this emission can contribute to the deviation of the braking index from 3, which is the standard prediction of the magnetic dipole radiation and aligned rotator mechanisms. A similar picture arises if one focuses on the second braking index. Moreover the braking index depends in a specific way on both the rotational frequency and the strength of the magnetic field which could facilitate the identification of this mechanism with respect to others. We also show that gravitational waves can be produced by polar gap regions due to their rapid charge-discharge process that takes place in timescales from nanosec to microsec. This can provide an alternative way to probe magnetars and test the polar gap model.
@@
Since the first detection of gravitational waves (GW) a few years ago LIGOScientific:2016aoc , a lot of effort has been concentrated on this front both theoretically and experimentally. With the advent of Pulsar Timing Arrays and a series of new interferometers that have been proposed and expected to enter operation in the near or far future, there has been a lot of interest for new astrophysical GW sources that can potentially reveal a lot of information regarding astrophysical and cosmological processes. One field that has been studied extensively for obvious reasons is the production of GW by neutron stars (NS). Due to their compactness, there is a lot of opportunity to emit GW under certain conditions. NS can get involved in binaries among themselves or with black holes. This is the strongest type of signal that can come from a NS and this has already been detected LIGOScientific:2017vwq . Other possible cases of GW emission by NS is via several non-radial oscillations modes that can be excited in principle either by internal processes or by encounters with bypassing objects. A rotating NS is more oblate in the equator. This by itself cannot produce any GW, since the latter requires a time varying quadrupole (or higher) moment. Therefore in case for example a bypassing object disturbs the rotation of the NS, a nonzero second derivative of the quadrupole moment can be created, leading to emission of GW. For gravitational waves produced by pulsating NS, see e.g. Andersson:2000mf . Another interesting possibility that could lead to an emission of a continuous GW signal is the existence of a thermal and/or magnetic mountain (see Gittins:2024zbg for a review). Thermal mountains can arise in situations where the NS accretes matter from a companion which compresses underlying layers, triggering nuclear reactions that alter locally the energy density and thus creating a quadrupole moment Ushomirsky:2000ax . Magnetic mountains can appear and produce GW in cases where there is a misalignment between magnetic and rotational axis. NS with large toroidal magnetic fields distort the NS into a prolate shape that can cause continuous GW emission Cutler:2002nw ; Glampedakis:2012qp .
@@
There is yet another subtle mechanism where GW can be produced by NS involving neither some internal process nor a bypassing object. As it is well known since the 60’s, a rotating NS with a magnetic field cannot be without the presence of a magnetosphere GJ . In case where the rotation and magnetic axes do not coincide, the co-rotating magnetosphere can create GW that in principle can be detected in future experiments. This scenario has been already studied e.g. in Nazari and Contopoulos where the GW produced by the magnetosphere were deduced. The focus of these two papers was on the detectability of these GW in upcoming interferometers. Continuous GW produced in the context of glitches has been considered in Yim:2023nda .
@@
In this paper we verify that indeed GW produced by the magnetosphere of NS with strong magnetic fields are within reach in upcoming experiments. In addition we show that this is a source of continuous GW with a spectrum consisting of two frequencies, i.e., the frequency of the NS rotation $f_{0}$ and $2f_{0}$ with a very specific relative amplitude between the two. Furthermore we focus on another aspect of this scenario. We study how the emission of such GW by the magnetosphere affects the first and second braking index of the NS. We shall argue that in case of magnetars with extremely high magnetic fields, the braking index can deviate from 3 which is the number expected in standard scenarios where the NS deceleration takes place due to magnetic dipole radiation and/or via torques of escaping to infinity currents from the poles of the NS. We find that unlike the usual magnetic dipole and aligned rotator models where the braking index is independent of both the rotation frequency and the magnetic field, in high magnetic field magnetars, the braking index depends on both mentioned quantities. Similar conclusions can be drawn for the second braking index. We will show that small deviations from the standard values result from NS with fields of at least $10^{15}-10^{16}$ Gauss. Although such high magnetic fields have not been observed in magnetars so far, this is not unexpected. Arguments based on the virial theorem suggest fields up to $10^{18}$ Gauss for NS with nuclear cores ShapiroLai , while for magnetars with a quark core this value can rise to $10^{20}$ Gauss Ferrer:2010wz . Hybrid stars with quark and nuclear matter and high magnetic fields have been studied in Sotani:2014rva . The effects of extremely high magnetic fields on the QCD equation of state (relevant also for NS) has been recently readdressed Fraga:2023cef ; Fraga:2023lzn .
@@ I Gravitational Waves produced by the Magnetosphere
[…]
@@
Figure 2: Upper panels: Deviation of the first braking index from the standard value 3 for magnetars of different magnetic field as a function of Log P (P being the period of the star in sec). Upper Left panel: From higher index value to lower, we have stars with $B=10^{16}$ and $5\times 10^{15}$ Gauss. Right panel: The same thing for $B=5\times 10^{17}$, $10^{17}$ and $5\times 10^{16}$ Gauss. Lower panels: Same as the upper panels for the deviation of the second braking index from the standard value 15. In all plots we use $R=20$ km and $\alpha=\pi/2$.
@@ II Braking Index
For any mechanism where $\dot{\omega}=-k\omega^{n}$ with $k$ independent of time, the braking index is n and the second braking index is m=n(2n1)m=n(2n-1). By inspection of Eq. (13) both the aligned rotator model and the magnetic dipole model give $n=3$ and $m=15$. Note also that GW emission provides first and second braking indices $n=5$ and $m=45$ respectively. The co-existence of the electromagnetic emission mechanisms with the GW emission one will make both $n$ and $m$ depart from their corresponding values of 3 and 15. This is depicted in Fig. (2) where both the first and second braking indices are plotted as a function of the rotational period. As it can be seen larger magnetic fields cause larger deviations from the standard value $n=3$. Starting from $\sim 10^{15}$ Gauss the deviation is $10^{-6}$, culminating to deviations of the order of $10^{-3}$ for fields of the order of $\sim 10^{17}$ Gauss. Such accuracy can be anticipated for some pulsars in the future given the current one which can be as low as $10^{-3}$. As expected, faster rotations cause in general larger deviations from the standard $n=3$ and $m=15$ values. A similar picture is drawn for the second braking index $m$. Note here that for a large part of the studied parameter space, it appears that the deviation from the standard value of $m$ is roughly an order of magnitude larger than that of $n$.
@@
Note also that although in general the deviation for both $n$ and $m$ increases as the period of the rotation decreases, this does happen indefinitely. As we approach msec periods both braking indices do not increase further but rather they start to decrease. This is more pronounced for larger magnetic fields. This is in distinct contrast to what happens to the braking index that is dominated by GW emission under the condition that the GW emission is due to misalignment between the principal axes and the rotation one, which is expected to have braking indices monotonically decreasing with the period. This feature is attributed to the nontrivial way that powers of $\omega R$ enter the expression of the quadrupole moment. It reflects the fact that the GW emission depends on both the rotation and magnetic field. Ordinarily, higher frequency means larger GW amplitude. However in this case, higher frequency leads also to smaller light cylinder and thus smaller volume for the magnetosphere that contributes to the GW emission. Another characteristic feature which is distinctly different from the typical GW emission due to the precession of a thermal mountain is the fact that in our studied scenario, both indices depend strongly on the magnetic field unlike the former case. Although the distinction of the presented mechanism with respect to GW emission from thermal mountains can be achieved by looking at isolated and non-accreting NS, the situation is more perplexed in the presence of magnetic mountains already mentioned in the introduction. Emission from this mechanism is also continuous in nature, it depends also on the magnetic field, and can dominate the emission from the magnetosphere. In principle both mechanisms are in place and both contributions should be taken into account. The magnetosphere contribution can be more pronounced if the NS does not possess large toroidal magnetic fields, condition that minimizes the contribution of the magnetic mountain. Therefore precise knowledge of the NS’s magnetic field is important. The magnetic field of NS is typically estimated by the deceleration rate of the rotation. However in principle, a different independent estimate of the magnetic field could actually test our scenario. In any case, even if the magnetic field cannot be determined independently, knowledge of both $n$ and $m$ can potentially distinguish this deceleration mechanism from the standard ones (magnetic dipole and aligned rotator) that are independent of the magnetic field since in our studied scenario both $n$ and $m$ depend on the magnetic field nontrivially. It could also potentially distinguish this mechanism from the magnetic mountain one, based on the fact that as the rotational frequency increases beyond a certain point, the former (in contrast to the latter) switches from an increasing GW power to a decreasing one. This is directly related to the aforementioned contraction of the light cylinder that beyond a point results to a reduction in the GW emission of the magnetosphere. The study of different NS with different rotational frequencies and magnetic fields could potentially disentangle the two contributions provided accurate measurements of the toroidal components of the magnetic fields.
As a final comment it should be noted that the inclination angle $\alpha$ and the magnetic fields evolve with time. In such a context the angle $\alpha$ and the magnetic field can be correlated to the rotational frequency. Understanding that correlation could provide another way to test this mechanism since it could make a concrete prediction on braking indices, magnetic field and $\alpha$ simultaneously. However for the moment this is an open problem. Both the time evolution of $\alpha$ and the magnetic field are not well understood basically because we do not know the exact mechanism of magnetic field formation in the star. In particular the time evolution of $\alpha$ can in principle change the values of the braking indices by changing the initial value of $\alpha$ and by contributing an extra term on $\dot{\omega}$ proportional to $\dot{\alpha}$. However the situation is far from clear and has not yet been resolved. Observations of several pulsars Weltevrede:2008ic indicate that there is a correlation between the period (and age) of the pulsar and the angle $\alpha$. Older pulsars seem to be more aligned than younger ones. Note that this study did not include millisecond pulsars which is our target group in this paper. Some 3D simulations of the magnetosphere Philippov:2013aha predict a rapid exponential or milder power law decay of $\alpha$ (depending on the type of the magnetosphere) taking place in the characteristic slow down timescale $(\dot{\omega}/\omega)^{-1}$. In any case, even for slower rotating pulsars it seems hard to reconcile an exponential or power law decay of the inclination angle with the observed pulsars since significant angles persist at ages where the simulations of Philippov:2013aha predict negligible ones (see e.g. the discussion in Johnston:2017wgm ). Even more puzzling is the fact that the Crab pulsar Lyne:2013voa ; Lyne:2014qqa exhibits an increase in the alignment angle with a rate of $0.566^{0}/$year contrary to the expectation of a decreasing $\alpha$ with time. Why $\alpha$ increases for the Crab pulsar and why this happens with that rate is not well established. The inclination angle could increase up to the time where dissipative torques due to viscosity kick in Jones . For that to happen, the star has to cool significantly for $10^{4}$ or $10^{5}$ years. The crucial element here is that this timescale is related to the cooling process and not to the slowing down of the star. As it is clear from the above discussion, the time evolution of the inclination angle is far from being understood. The purpose of this paper here is not to speculate on these open issues, but rather to demonstrate that under general and robust conditions, the braking index due to GW emission can deviate significantly from the predicted values of the standard scenarios of the magnetic dipole radiation and the aligned rotator. Fig. 2 provides braking indices for an angle $\alpha=\pi/2$. Because of the fact that the $\alpha$ dependence of the function $f_{E}$ enters as powers of $\sin 2\alpha$ and $\sin 4\alpha$, the derivative of $\alpha$ with respect to time is zero at that exact value and therefore does not contribute an extra term in the braking index. Therefore in the specific case the results are exact (i.e. the effect of a nonzero $\dot{\alpha}$ is taken into account). As a further cross-check the braking index for a different value e.g., $\alpha=\pi/4$ has been calculated, using a rate of change for $\alpha$ similar to the observed one from the Crab pulsar which has several of the desired properties (large magnetic field $\sim 10^{13}$ Gauss field and a small 33 msec period). Including the effect of the time evolution of inclination angle is negligible (changes of the order of $10^{-16}$). In fact this does not come as a surprise. The functional dependence of $f_{E}$ is power law for both $\omega R$ and $\sin\alpha$. Therefore the relative strength of the extra term that involves $\dot{\alpha}$ comes down to the comparison between the timescales $(\dot{\alpha}/\alpha)^{-1}$ and $(\dot{\omega}/\omega)^{-1}$. As long as the timescale for the decay of $\alpha$ is much longer than that of the rotational deceleration, the effect will be negligible. Obviously if the timescales become comparable, the contributions become comparable too. However even in that extreme case, it is highly unlike to have the two contributions fine tuned in a way that one cancels the other. Therefore one anticipates that the values presented in Fig. 2 even in that special case give the correct ballpark of the braking index. The purpose of this work is to point out that GW emission from the magnetosphere can cause observational changes in the braking index of the magnetars with a nontrivial dependence on the magnetic field in striking contrast to the aligned rotator and the magnetic dipole models as well as the emission of GW due to oblateness. As long as $(\dot{\alpha}/\alpha)^{-1}>>(\dot{\omega}/\omega)^{-1}$ we can ignore the time dependence of $\alpha$ since it is negligible. We leave for future work the interesting case where the two timescale are comparable.
@@ III Gravitational Waves from the Polar Cap
where $z$ denotes the coordinate along the rotation axis. Note that during the charging of the gap t=zRT/2t=z-R-T/2 for $2\leq t\leq-T/2+h. TT$ is the total period of the cycle charge-discharge that we will estimate later on. Within our approximation after the electric field reaches its maximum, it will stay there up to time t=T/2ht=T/2-h where the spark ignites. After that moment the charge redistributes and reaches the Goldreich-Julian value that screens completely the electric field. Therefore after t=T/2ht=T/2-h and until $T=T/2$, we have t=T/2+Rzt=T/2+R-z. Note also that $z$ is measured from the center of mass which in good approximation is taken to be the center of the star. Fig. (3) illustrates the geometry of the polar cap region.
@@
By inspection of Eq. (20), for relatively slow rotations $\ddot{M}_{33}>>\ddot{M}_{11}$ because $R>>r_{p}\simeq R\sqrt{\omega R}$. To give an approximate estimate for the GW amplitude, we note that $g(t)$ is of order one for time between $T/2\leq t\leq-T/2+h$ and $T/2-h\leq t\leq T/2$. As we mentioned earlier the Fourier analysis gives a peak at $f_{1}=h^{-1}$ and a smaller amplitude at $f_{0}=T^{-1}$, consistent with our argument here.
@@
We can give now a rough estimate of the timescale needed to produce enough charge in the gap vacuum to discharge it. Within the timescale of $h$ which is the time it takes for a particle to cross the gap, $N_{\gamma}$ curvature photons are produced, which at time $2h$ each of them has produced one electron-positron pair. The amount of required charge in order to discharge the gap is given by the Goldreich-Julian density which over the poles is $\rho_{e}=-\vec{\omega}\cdot\vec{B}/(2\pi)$. Therefore the amount of charge needed is $|Q|=\rho_{e}\pi r_{p}^{2}h$. Within this simple argument, the number of pairs grows as a geometrical series every $2h$, leading to $T\sim 2h\ln[(2N_{\gamma}-1)Q/(2N_{\gamma}e)]/\ln(2N_{\gamma})$. For example for a NS with $B_{12}=P=\rho_{6}=1$, we get $T\sim 24h$, while for $B_{12}=\rho_{6}=1$ and $P=10^{-2}$, we get $T\sim 38$. These results are comparable to the range given in Ruderman . However we should stress that in order to discharge the gap, two conditions have to be fulfilled: i) the mean free path for converting the curvature photons to electron-positron pairs has to be smaller than $h$ and ii) $N_{\gamma}$ must be larger than 0.5 or more in order for the pair creation to grow at each time $h$. $h$ cannot become larger than $r_{p}$ and therefore as a NS slows down and $\omega$ decreases, the gap increases. When the gap becomes $h\sim r_{p}$, the pulsar ceases to emit because there can be no discharge and production of electron-positron pairs. We are interested in NS with large magnetic fields. As the magnetic field increases, Eq. (22) shows that $h$ drops and subsequently $N_{\gamma}$ drops too (see Eq. (24)). Therefore at some magnetic field, $N_{\gamma}$ becomes smaller than 0.5 and there is no cascade of electron-positron pairs created. In this case it is expected that $h$ will increase as much needed, as long as $h\leq r_{p}$, until $N_{\gamma}$ becomes of order 1.
@@
As we mentioned $\ddot{M}_{11}<<\ddot{M}_{33}$. The above equation gives
$h_{+}=2\times 10^{-29}P_{3}^{-3}B_{15}^{2}h_{10}r_{1}^{-1}\sin^{2}\psi,$ (26)
where $P_{3}$ is the period of rotation in units of $10^{-3}$ sec, $B_{15}$ the magnetic field on the poles in units of $10^{15}$ Gauss, $h_{10}$ the gap in units of 10 cm and $r_{1}$ the distance to the NS in kpc. For this formula we have assumed $R=20$ km and $\rho_{6}=1$. Recall that $h$ is not a free parameter but it is the maximum between the value of Eq. (22) and the minimum value required in order to have $N_{\gamma}\sim 1$, so the discharge process can take place. Ignoring the factor $\sin^{2}\psi$ and setting the distance to the source to 1kpc, we can estimate the amplitude for different benchmark points presented in Table 1. $f_{0}$ and $f_{1}$ are the two frequencies mentioned earlier. The amplitude for $h_{+}$ that is presented refers to $f_{1}$. Depending on the exact form of the functions $f(t)$ or $g(t)$, the amplitude at $f_{0}$ will be somewhat suppressed compared to that at $f_{1}$ that we quote at the table. For all benchmark points we used a period for the magnetar of $10^{-3}$ sec. Although for the moment there are no current experiments probing the MHz-GHz range of frequencies, several proposals for future experiments exist, aiming exactly at detection of high frequency GW Aggarwal:2020olq . Table 1 of Aggarwal:2020olq gives a summary of all the high GW frequency proposals. Among the different proposals, detectors based on enhanced magnetic conversion could in principle operate at $\sim 10$ GHz with a sensitivity for the strain $10^{-30}$ to $10^{-26}$. Similarly for superconducting rings, the projected sensitivity at 10 GHz is $10^{-31}$. By inspection of Table 1 magnetic fields as low as $10^{14}$ Gauss are above the threshold in the GHz frequency range of the aforementioned proposals.